# Quantum information scrambling in adiabatically-driven critical systems

### v1

@@ Abstract
Quantum information scrambling refers to the spread of the initially stored information over many degrees of freedom of a quantum many-body system. Information scrambling is intimately linked to the thermalization of isolated quantum many-body systems and has been typically studied in a sudden quench scenario. Here we extend the notion of quantum information scrambling to critical quantum many-body systems undergoing an adiabatic evolution. In particular, we analyze how the symmetry-breaking information of an initial state is scrambled in adiabatically-driven integrable systems, such as the Lipkin-Meshkov-Glick and quantum Rabi model. Following a time-dependent protocol that drives the system from a symmetry-breaking to a normal phase, we show how the initial information is scrambled even for perfect adiabatic evolutions as indicated by the expectation value of a suitable observable. We detail the underlying mechanism for quantum information scrambling, its relation to ground- and excited-state quantum phase transitions and quantify the degree of scrambling in terms of the number of eigenstates that participate in the encoding of the initial symmetry-breaking information. While the energy of the final state remains unaltered in an adiabatic protocol, the relative phases among eigenstates are scrambled and so is the symmetry-breaking information. We show that a potential information retrieval, following a time-reversed protocol, is hindered by small perturbations as indicated by a vanishingly small Loschmidt echo and out-of-time-ordered correlators. The reported phenomenon is amenable for its experimental verification and may help in the understanding of information scrambling in critical quantum many-body systems.
Keywords: Nonequilibrium critical dynamics, quantum phase transitions, quantum information scrambling
@@ 1 Introduction
The process by which quantum information, initially localized within the degrees of freedom of a many-body system, propagates and becomes distributed across the entire system degrees of freedom is known as information scrambling [1, 2, 3]. This phenomenon is particularly intriguing from a fundamental perspective, as it delves into the mechanisms of information dispersal and entanglement in complex quantum systems [4, 5, 6]. Understanding how this scrambling occurs can shed light on the underlying principles governing quantum dynamics and may have significant implications in quantum computing [7], quantum chaos [8], quantum thermodynamics [9, 10, 11, 6], and high energy physics [12].
@@
Quantum many-body systems offer an unrivaled testbed platform to explore the rich interplay between entanglement, thermalization, non-equilibrium dynamics, and even quantum phase transitions (QPTs) [13, 14, 15]. Quantum information scrambling has been primarily linked with thermalization in closed systems, and thus it is intimately related to chaotic behavior in quantum many-body systems [13, 16, 17, 18, 9, 19, 20, 21]. The complex nature of correlations in non-integrable systems reveals itself in signatures of information scrambling [22], typically measured in terms of out-of-time-ordered correlation (OTOCs) functions and the Loschmidt echo [1, 2, 3]. Such scrambling of information has been extensively studied in a sudden-quench scenario [23, 24, 25, 26, 19, 20, 27, 28], where an initial state evolves under its governing and constant Hamiltonian. The integrability of the system has an profound impact on how the information gets scrambled, being non-integrable systems better scramblers as evidenced by the dynamics of OTOCs and Loschmidt echo (see for example [26]). The role of time-dependent quenches in quantum information scrambling remains, however, largely unexplored, and more so when involving critical features of a quantum many-body system.
@@
In this work, we aim at closing this gap and analyze quantum information scrambling in critical systems undergoing a fully coherent adiabatic protocol. The considered systems feature a normal and symmetry-breaking phase where a $\mathbb{Z}_{2}$ parity is spontaneously broken. The initial information consists in one of the two possible symmetry-breaking choices and is encoded in an initial state that populates a large collection of degenerate energy eigenstates that break the discrete parity symmetry of the governing Hamiltonian. Such symmetry-breaking information remains unaltered under the dynamics of the initial Hamiltonian. This is possible for systems where a symmetry-breaking phase is not only restricted to the ground and first excited states, but rather it propagates up to a certain critical excitation energy [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. This is indeed the case for many-body systems with few effective and collective degrees of freedom, such as the Lipkin-Meshkov-Glick (LMG) model [43, 44, 45, 46, 47, 48], Dicke model [49, 50] or quantum Rabi model (QRM) [51, 52, 53, 54, 55]. These systems feature a ground-state mean-field quantum phase transition, as well as an excited-state quantum phase transition precisely at a critical excitation energy that divides the spectrum between normal (energy eigenstates with well-defined parity) and a degenerate phase, where eigenstates with opposite parities are degenerated by pairs, as described by a standard double-well semiclassical potential. We stress that the LMG and the QRM are integrable systems, while the Dicke model exhibits a region with chaotic behavior at sufficiently high excitation energies [56, 57, 58, 59]. It is worth remarking that the paradigmatic one-dimensional Ising model with a transverse field does not meet the previous criterion since only the two lowest energy eigenstates are degenerate in the antiferro- or ferromagnetic phase separated by the quantum phase transition [15].
@@
The adiabatic protocol brings the initial symmetry-breaking state towards the normal and back to the degenerate phase by quenching the control parameter of the system. Albeit the populations of the energy eigenstates remain constant under the adiabatic theorem, the symmetry-breaking information is scrambled due to the different and non-commensurable relative phases gained during the protocol, provided the system is driven into the normal phase. Such dynamical phases are uniformly distributed resulting in an adiabatic quantum information scrambling (AQIS). The effectiveness of AQIS is illustrated in the LMG model, and quantified in terms of the final expectation value of a suitable symmetry-breaking observable, the number of populated eigenstates, as well as with the OTOCs. The robustness of the AQIS is analyzed by means of the Loschmidt echo between an adiabatically evolved state, and its time-reversed protocol with a small perturbation. The AQIS is illustrated for different initial states and another integrable quantum critical model, namely, the QRM [51].
@@
The article is organized as follows. In Sec. 2 we present the mechanism that leads to the quantum scrambling of symmetry-breaking information for an adiabatically-driven quantum many-body system, i.e. the AQIS. In Sec. 3 we first introduce the LMG model and then present numerical results of the adiabatic quantum information scrambling. A different model, the QRM, is analyzed in Sec. 4 and presents similar numerical results supporting the quantum information scrambling. Finally, in Sec. 5 we summarize the main results of the article.
@@ 2 Adiabatic Quantum Information Scrambling
Let us start considering a quantum critical system described by a Hamiltonian $\hat{H}(g)$ that depends on an external controllable parameter $g$, such that
@@
where $E_{k}(g)$ denote the energy of the $k$th eigenstate $\left|\phi_{k}(g)\rangle\right.$ for $g$. In addition, we assume that the Hamiltonian commutes with a discrete parity operator $\hat{\Pi}$, $[\hat{H}(g),\hat{\Pi}]=0$ so that the eigenstates can be labeled in terms of the eigenvalues of $\hat{\Pi}$, $+1$ or $-1$, owing to a $\mathbb{Z}_{2}$ symmetry, i.e. the eigenstates can be written as $\left|\phi_{k,\pm}(g)\rangle\right.$ with $\hat{\Pi}\left|\phi_{k,\pm}(g)\rangle\right.=\pm 1\left|\phi_{k,\pm}(g)\rangle\right.$, and similarly for its energy $E_{k,\pm}(g)$, so that
@@
The phase diagram of the Hamiltonian can be split in two regions, namely, normal phase where $E_{k,+}(g)\neq E_{k,-}(g)$ and a symmetry-breaking phase where the eigenstates belonging to a different parity subspace are degenerate, i.e. $E_{k,+}(g)=E_{k,-}(g)$, possibly up to a certain critical excitation energy $E_{c}(g)$. These two phases are separated by ground- and excited-state quantum phase transition [15, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], taking place a critical value $g_{c}$ and at an excitation energy $E_{c}(g)$, respectively. Note that in the symmetry-breaking phase, any state of the form $\left|\phi(g)\rangle\right.=\alpha\left|\phi_{k,+}(g)\rangle\right.+\beta\left|\phi_{k,-}(g)\rangle\right.$ with $|\alpha|^{2}+|\beta|^{2}=1$ and $E_{k,+}(g)=E_{k,-}(g)$ is also a valid eigenstate of $\hat{H}(g)$. Yet, for any $\alpha$ such that $|\alpha|\neq 1,0$, $\left|\phi(g)\rangle\right.$ is no longer an eigenstate of the parity operator, $\hat{\Pi}\left|\phi(g)\rangle\right.\neq\pm\left|\phi(g)\rangle\right.$, and thus the symmetry may be spontaneously broken. This effect can be quantified employing the expectation value of a suitable operator $\hat{\mathcal{O}}$, such that $\left.\langle\phi_{k,+}(g_{0})\right|\hat{\mathcal{O}}\left|\phi_{k,-}(g_{0})\rangle\right.\neq 0$ and $\left.\langle\phi_{k,p}(g_{0})\right|\hat{\mathcal{O}}\left|\phi_{k,p}(g_{0})\rangle\right.=0$. Indeed, its expectation value behaves as an order parameter, i.e. $\langle\hat{\mathcal{O}}\rangle=0$ for a symmetric state and $\langle\hat{\mathcal{O}}\rangle\neq 0$ for symmetry-breaking states, as it is the case for the magnetization in a standard ferro-to-paramagnetic phase transition.
We consider an initial state $\left|\psi_{0}\rangle\right.$ which breaks this symmetry for an initial value $g_{0}$ of the control parameter in the symmetry-breaking phase, i.e. $\left.\langle\psi_{0}\right|\hat{\mathcal{O}}\left|\psi_{0}\rangle\right.\neq 0$. This is the information we will consider throughout the rest of the article. Note that in general $[\hat{H}(g),\hat{\mathcal{O}}]\neq 0$. Such state can be written in the eigenbasis of $\hat{H}(g_{0})$ as
@@
The coefficients $c_{k,p}$ encode therefore the choice of the symmetry-breaking state, and thus the initial information. Moreover, since $\hat{\mathcal{O}}=\sum_{n}\gamma_{n}\left|\gamma_{n}\rangle\right.\left.\langle\gamma_{n}\right|$, we can characterize the probability distribution of measuring $\gamma_{n}$’s in the initial state $\left|\psi_{0}\rangle\right.$, which reads as
@@
For a symmetric state, $P(\gamma_{n})=P(-\gamma_{n})$ and thus $\langle\hat{\mathcal{O}}\rangle=0$. However, for maximally symmetry-broken states, the distribution is only non-zero for one of the two branches, i.e. either $P(\gamma>0)\neq$ or $P(\gamma<0)\neq 0$. For simplicity, and without loss of generality, we will consider initial maximally symmetry-broken states $\left|\psi_{0}\rangle\right.$ where $P(\gamma>0)\neq 0$ and $P(\gamma<0)=0$. In addition, note that the energy probability distribution of the initial state is simply given by $P(E_{k}(g))=|c_{k,+}|^{2}+|c_{k,-}|^{2}$, since we consider the initial state to be in the symmetry-breaking phase. Therefore, $P(E)$ is independent of the initial information, and thus the energy probability distribution does not depend on how the symmetry is broken.
We then consider a time-dependent protocol $g(t)$ that drives the system from $g(0)=g_{0}$ to $g(\tau)=g_{1}$, and back to the initial control value, $g(2\tau)=g_{0}$. For simplicity, we consider a linear ramp, namely,
@@
where $\varphi_{k,p}$ denotes the accumulated phase for the $k$-th eigenstate of parity $p$ upon completing the cycle from $g_{0}$ to $g_{1}$ and back to $g_{0}$, i.e.
@@
Under the adiabatic approximation, neither the average energy nor the energy probability distribution $P(E)$ are altered. The relative phases between degenerate eigenstates are irrelevant to the energy, as commented above. However, the symmetry-breaking information may be largely affected since $\left.\langle\phi_{k,+}(g_{0})\right|\hat{\mathcal{O}}\left|\phi_{k,-}(g_{0})\rangle\right.\neq 0$, and thus $\left.\langle\psi(2\tau)\right|\hat{\mathcal{O}}\left|\psi(2\tau)\rangle\right.$ will depend on the phases $\varphi_{k,p}$, and will be, in general, different from the initial condition $\left.\langle\psi(2\tau)\right|\hat{\mathcal{O}}\left|\psi(2\tau)\rangle\right.\neq\left.\langle\psi_{0}\right|\hat{\mathcal{O}}\left|\psi_{0}\rangle\right.$.
@@
Now, since only the eigenstates with the same $k$ label and opposite parity are degenerate, then we can further simplify the previous expression and find
@@
regardless of the specific time $\tau$. Note that the phase differences $\delta\varphi_{k}$ will only be non-zero if the system is driven from the symmetry-breaking to the normal phase, as otherwise $E_{k,+}=E_{k,-}$ leading to $\delta\varphi_{k}=0$. Hence, the mechanism for AQIS requires $g_{0}$ to be in the symmetry-breaking and $g_{1}$ in the normal phase, thus forcing the system to traverse the excited-state quantum phase transition.
A consequence of Eq. (10) is that the probability distribution of the operator $\hat{\mathcal{O}}$ for the final state $\left|\psi(2\tau)\rangle\right.$ will exhibit a balanced distribution for $\gamma<0$ and $\gamma>0$, as it were a symmetric state. In this manner, the initial symmetry-breaking information is scrambled in the quantum system due to the large number of populated eigenstates. We can anticipate that the expression given in Eq. (10), i.e. AQIS will be more effective the larger the support of the initial state over the eigenstates of $\hat{H}(g_{0})$. In the following, we illustrate the AQIS in the LMG and QRM models, supporting with numerical simulations its effectiveness, robustness, and dependence on different initial states.
3 Lipkin-Meshkov-Glick model
The LMG model [43], originally introduced in the context of nuclear physics, describes the long-range dipole-dipole interaction of $N$ spin-$\frac{1}{2}$ under a transverse magnetic field. This model exhibits a QPT [15] at a certain critical value of the field strength $g_{c}$ [44, 45, 46, 47, 48], and has attracted renewed attention due to its experimental realization, e.g. with cold atoms [60] or in trapped ions [61]. Indeed, the LMG has served as a test bed for the exploration of different aspects of quantum critical systems [62, 63, 64, 65, 66, 67, 37, 68, 69, 70, 71, 72, 73, 74, 75, 39, 76]. The Hamiltonian of the $N$ spin-$\frac{1}{2}$ particles can be written as (unit of frequency equal to $1$, and $\hbar=1$)

### v2

@@ Abstract
Quantum information scrambling refers to the spread of the initially stored information over many degrees of freedom of a quantum many-body system. Information scrambling is intimately linked to the thermalization of isolated quantum many-body systems, and has been typically studied in a sudden quench scenario. Here, we extend the notion of quantum information scrambling to critical quantum many-body systems undergoing an adiabatic evolution. In particular, we analyze how the symmetry-breaking information of an initial state is scrambled in adiabatically driven integrable systems, such as the Lipkin–Meshkov–Glick and quantum Rabi models. Following a time-dependent protocol that drives the system from symmetry-breaking to a normal phase, we show how the initial information is scrambled, even for perfect adiabatic evolutions, as indicated by the expectation value of a suitable observable. We detail the underlying mechanism for quantum information scrambling, its relation to ground- and excited-state quantum phase transitions, and quantify the degree of scrambling in terms of the number of eigenstates that participate in the encoding of the initial symmetry-breaking information. While the energy of the final state remains unaltered in an adiabatic protocol, the relative phases among eigenstates are scrambled, and so is the symmetry-breaking information. We show that a potential information retrieval, following a time-reversed protocol, is hindered by small perturbations, as indicated by a vanishingly small Loschmidt echo and out-of-time-ordered correlators. The reported phenomenon is amenable for its experimental verification, and may help in the understanding of information scrambling in critical quantum many-body systems.
Keywords: nonequilibrium critical dynamics; quantum phase transitions; quantum information scrambling
@@ 1 Introduction
The process by which quantum information, initially localized within the degrees of freedom of a many-body system, propagates and becomes distributed across the entire system degrees of freedom, is known as information scrambling [1, 2, 3]. This phenomenon is particularly intriguing from a fundamental perspective, as it delves into the mechanisms of information dispersal and entanglement in complex quantum systems [4, 5, 6]. Understanding how this scrambling occurs can shed light on the underlying principles governing quantum dynamics, and may have significant implications in quantum computing [7], quantum chaos [8], quantum thermodynamics [9, 10, 11, 6], and high energy physics [12].
@@
Quantum many-body systems offer an unrivaled testbed platform to explore the rich interplay between entanglement, thermalization, non-equilibrium dynamics, and even quantum phase transitions (QPTs) [13, 14, 15]. Quantum information scrambling has been primarily linked with thermalization in closed systems and, thus, it is intimately related to chaotic behavior in quantum many-body systems [13, 16, 17, 18, 9, 19, 20, 21]. The complex nature of correlations in non-integrable systems reveals itself in signatures of information scrambling [22], typically measured in terms of out-of-time-ordered correlation (OTOC) functions [23], tripartite mutual information [24, 25], and the Loschmidt echo [1, 2, 3]. Such scrambling of information has been extensively studied in a sudden-quench scenario [26, 27, 28, 29, 19, 20, 30, 31], where an initial state evolves under its governing and constant Hamiltonian. The integrability of the system has a profound impact on how the information becomes scrambled, with non-integrable systems being better scramblers, as evidenced by the dynamics of OTOCs and the Loschmidt echo (see, for example, [29]). Additionally, OTOCs can be utilized to dynamically detect both equilibrium and nonequilibrium phase transitions in various many-body quantum systems [32, 21, 33, 34, 35]. However, the role of time-dependent quenches in quantum information scrambling remains largely unexplored, and even more so when involving critical features of a quantum many-body system.
@@
In this work, we aim to close this gap and analyze quantum information scrambling in critical systems undergoing a fully coherent adiabatic protocol. The adiabatic protocol ensures zero irreversible work, which may otherwise be significant in critical systems [36, 37, 38, 39, 40]. The considered systems feature a normal and symmetry-breaking phase where a $\mathbb{Z}_{2}$ parity is spontaneously broken. The initial information consists of one of the two possible symmetry-breaking choices, and is encoded in an initial state that populates a large collection of degenerate energy eigenstates that break the discrete parity symmetry of the governing Hamiltonian. Such symmetry-breaking information remains unaltered under the dynamics of the initial Hamiltonian. This is possible for systems where a symmetry-breaking phase is not only restricted to the ground and first excited states, but rather propagates up to a certain critical excitation energy [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Indeed, this is the case for many-body systems with few effective and collective degrees of freedom, such as the Lipkin–Meshkov–Glick (LMG) model [55, 56, 57, 58, 59, 60], Dicke model [61, 62], or quantum Rabi model (QRM) [63, 64, 65, 66, 67]. These systems feature a ground-state mean-field quantum phase transition, as well as an excited-state quantum phase transition precisely at a critical excitation energy that divides the spectrum between normal (energy eigenstates with well-defined parity) and a degenerate phase, where eigenstates with opposite parities are degenerated by pairs, as described by a standard double-well semiclassical potential. We stress that the LMG and QRM are integrable systems, while the Dicke model exhibits a region with chaotic behavior at sufficiently high excitation energies [68, 69, 70, 71]. It is worth remarking that the paradigmatic one-dimensional Ising model with a transverse field does not meet the previous criterion, since only the two lowest-energy eigenstates are degenerate in the antiferro- or ferromagnetic phase separated by the quantum phase transition [15].
@@
The adiabatic protocol brings the initial symmetry-breaking state towards the normal and back to the degenerate phase by quenching the control parameter of the system. However, while the populations of the energy eigenstates remain constant under the adiabatic theorem, the symmetry-breaking information is scrambled, due to the different and non-commensurable relative phases gained during the protocol, provided that the system is driven into the normal phase. Such dynamical phases are uniformly distributed, resulting in adiabatic quantum information scrambling (AQIS). The effectiveness of AQIS is illustrated in the LMG model, and quantified in terms of the final expectation value of a suitable symmetry-breaking observables and the number of populated eigenstates, as well as with the OTOCs. The robustness of AQIS is analyzed by means of the Loschmidt echo between an adiabatically evolved state and its time-reversed protocol with a small perturbation. AQIS is illustrated for different initial states and another integrable quantum critical model, namely QRM [63].
@@
The article is organized as follows. In Section 2 we present the mechanism that leads to the quantum scrambling of symmetry-breaking information for an adiabatically driven quantum many-body system, i.e., AQIS. In Section 3 we first introduce the LMG model, and then present numerical results of the adiabatic quantum information scrambling. A different model, QRM, is analyzed in Section 4, and presents similar numerical results, supporting quantum information scrambling. Finally, in Section 5, we summarize the main results of the article.
@@ 2 Adiabatic Quantum Information Scrambling
Let us start by considering a quantum critical system described by a Hamiltonian $\hat{H}(g)$ that depends on an external controllable parameter $g$, such that
@@
where $E_{k}(g)$ denotes the energy of the $k$th eigenstate $\left|\phi_{k}(g)\rangle\right.$ for $g$. In addition, we assume that the Hamiltonian commutes with a discrete parity operator $\hat{\Pi}$, $[\hat{H}(g),\hat{\Pi}]=0$, so that the eigenstates can be labeled in terms of the eigenvalues of $\hat{\Pi}$, $+1$ or $-1$, owing to a $\mathbb{Z}_{2}$ symmetry, i.e., the eigenstates can be written as $\left|\phi_{k,\pm}(g)\rangle\right.$ with $\hat{\Pi}\left|\phi_{k,\pm}(g)\rangle\right.=\pm 1\left|\phi_{k,\pm}(g)\rangle\right.$, and similarly for its energy $E_{k,\pm}(g)$, so that
@@
The phase diagram of the Hamiltonian can be split in two regions, namely the normal phase where $E_{k,+}(g)\neq E_{k,-}(g)$, and a symmetry-breaking phase where the eigenstates belonging to a different parity subspace are degenerate, i.e., $E_{k,+}(g)=E_{k,-}(g)$, possibly up to a certain critical excitation energy $E_{c}(g)$. These two phases are separated by ground- and excited-state quantum phase transitions [15, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50], taking the place of a critical value $g_{c}$ and at an excitation energy $E_{c}(g)$, respectively. Note that in the symmetry-breaking phase, any state of the form $\left|\phi(g)\rangle\right.=\alpha\left|\phi_{k,+}(g)\rangle\right.+\beta\left|\phi_{k,-}(g)\rangle\right.$ with $|\alpha|^{2}+|\beta|^{2}=1$ and $E_{k,+}(g)=E_{k,-}(g)$ is also a valid eigenstate of $\hat{H}(g)$. However, for any $\alpha$ such that $|\alpha|\neq 1,0$, $\left|\phi(g)\rangle\right.$ is no longer an eigenstate of the parity operator, $\hat{\Pi}\left|\phi(g)\rangle\right.\neq\pm\left|\phi(g)\rangle\right.$, and thus the symmetry may be spontaneously broken. This effect can be quantified employing the expectation value of a suitable operator $\hat{\mathcal{O}}$, such that $\left.\langle\phi_{k,+}(g_{0})\right|\hat{\mathcal{O}}\left|\phi_{k,-}(g_{0})\rangle\right.\neq 0$ and $\left.\langle\phi_{k,p}(g_{0})\right|\hat{\mathcal{O}}\left|\phi_{k,p}(g_{0})\rangle\right.=0$. Indeed, its expectation value behaves as an order parameter, i.e., $\langle\hat{\mathcal{O}}\rangle=0$ for a symmetric state, and $\langle\hat{\mathcal{O}}\rangle\neq 0$ for symmetry-breaking states, as is the case for magnetization in a standard ferro-to-paramagnetic phase transition.
We consider an initial state, $\left|\psi_{0}\rangle\right.$, which breaks this symmetry for an initial value $g_{0}$ of the control parameter in the symmetry-breaking phase, i.e., $\left.\langle\psi_{0}\right|\hat{\mathcal{O}}\left|\psi_{0}\rangle\right.\neq 0$. This is the information we will consider throughout the rest of the article. Note that, in general, $[\hat{H}(g),\hat{\mathcal{O}}]\neq 0$. Such a state can be written in the eigenbasis of $\hat{H}(g_{0})$ as
@@
The coefficients, $c_{k,p}$, therefore encode the choice of the symmetry-breaking state, and thus the initial information. Moreover, since $\hat{\mathcal{O}}=\sum_{n}\gamma_{n}\left|\gamma_{n}\rangle\right.\left.\langle\gamma_{n}\right|$, we can characterize the probability distribution of measuring $\gamma_{n}$’s in the initial state $\left|\psi_{0}\rangle\right.$, which reads as
@@
For a symmetric state, $P(\gamma_{n})=P(-\gamma_{n})$, and thus $\langle\hat{\mathcal{O}}\rangle=0$. However, for maximally symmetry-broken states, the distribution is only non-zero for one of the two branches, i.e., either $P(\gamma>0)\neq$ or $P(\gamma<0)\neq 0$. For simplicity, and without loss of generality, we will consider initial maximally symmetry-broken states $\left|\psi_{0}\rangle\right.$ where $P(\gamma>0)\neq 0$ and $P(\gamma<0)=0$. In addition, note that the energy probability distribution of the initial state is simply given by $P(E_{k}(g))=|c_{k,+}|^{2}+|c_{k,-}|^{2}$, since we consider the initial state to be in the symmetry-breaking phase. Therefore, $P(E)$ is independent of the initial information, and thus the energy probability distribution does not depend on how the symmetry is broken.
We then consider a time-dependent protocol, $g(t)$, that drives the system from $g(0)=g_{0}$ to $g(\tau)=g_{1}$, and back to the initial control value, $g(2\tau)=g_{0}$. For simplicity, we consider a linear ramp, namely
@@
where $\varphi_{k,p}$ denotes the accumulated phase for the $k$-th eigenstate of parity $p$ upon completing the cycle from $g_{0}$ to $g_{1}$ and back to $g_{0}$, i.e.,
@@
Under the adiabatic approximation, neither the average energy nor the energy probability distribution $P(E)$ are altered. The relative phases between degenerate eigenstates are irrelevant to the energy, as commented above. However, the symmetry-breaking information may be largely affected, since $\left.\langle\phi_{k,+}(g_{0})\right|\hat{\mathcal{O}}\left|\phi_{k,-}(g_{0})\rangle\right.\neq 0$, and thus $\left.\langle\psi(2\tau)\right|\hat{\mathcal{O}}\left|\psi(2\tau)\rangle\right.$ will depend on the phases $\varphi_{k,p}$, and will be, in general, different from the initial condition $\left.\langle\psi(2\tau)\right|\hat{\mathcal{O}}\left|\psi(2\tau)\rangle\right.\neq\left.\langle\psi_{0}\right|\hat{\mathcal{O}}\left|\psi_{0}\rangle\right.$.
@@
Now, since only the eigenstates with the same $k$ label and opposite parity are degenerate, we can further simplify the previous expression, and find that
@@
regardless of the specific time $\tau$. Note that the phase differences $\delta\varphi_{k}$ will only be non-zero if the system is driven from the symmetry-breaking to the normal phase, as $E_{k,+}=E_{k,-}$ otherwise, leading to $\delta\varphi_{k}=0$. Hence, the mechanism for AQIS requires $g_{0}$ to be in the symmetry-breaking phase and $g_{1}$ in the normal phase, thus forcing the system to traverse the excited-state quantum phase transition.
A consequence of Equation (10) is that the probability distribution of the operator $\hat{\mathcal{O}}$ for the final state $\left|\psi(2\tau)\rangle\right.$ will exhibit a balanced distribution for $\gamma<0$ and $\gamma>0$, as if it were a symmetric state. In this manner, the initial symmetry-breaking information is scrambled in the quantum system due to the large number of populated eigenstates. We can anticipate the expression given in Equation (10), i.e., that AQIS will be more effective the larger the support of the initial state over the eigenstates of $\hat{H}(g_{0})$. In the following, we illustrate AQIS in the LMG and QRM models, supporting its effectiveness, robustness, and dependence on different initial states with numerical simulations.
3 Lipkin–Meshkov–Glick Model
The LMG model [55], originally introduced in the context of nuclear physics, describes the long-range dipole–dipole interaction of $N$ spin-$\frac{1}{2}$ under a transverse magnetic field. This model exhibits QPT [15] at a certain critical value of the field strength $g_{c}$ [56, 57, 58, 59, 60], and has attracted renewed attention due to its experimental realization, e.g., with cold atoms [72] or in trapped ions [73]. Indeed, the LMG has served as a test bed for the exploration of different aspects of quantum critical systems [74, 75, 76, 77, 78, 79, 49, 80, 81, 82, 83, 84, 85, 86, 39, 51, 87]. The Hamiltonian of the $N$ spin-$\frac{1}{2}$ particles can be written as (where the unit of frequency is equal to $1$, and $\hbar=1$)